OpenStudy (anonymous):

PLEASE HELP WILL GIVE A MEDAL Find the derivative of f(x) = 8x + 4 at x = 9.

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

Find the derivative first and then plug in x = 9. What is the derivative of f(x) = 8x+4

OpenStudy (anonymous):

It might be a bit tricky since you have x = 9, once you find the derivative, see what you can say about x.

OpenStudy (freckles):

this is just a line you could find the derivative easily if you recall some algebra from your algebra days

OpenStudy (freckles):

slope=derivative after all

OpenStudy (anonymous):

so if the derivative is 8 what do I do with the 9?

OpenStudy (freckles):

the derivative is 8 for all x

OpenStudy (anonymous):

how would I figure it out if the equation was quadratic?

OpenStudy (anonymous):

Find the derivative of f(x) = -4x2 + 11x at x = 10

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

\[f(x)= x^n~~~~\implies f'(x) = nx^{n-1}\] take the derivative term by term as previously.

OpenStudy (anonymous):

I still don't understand

OpenStudy (freckles):

to find the derivative of x^2 use the formula @iambatman has given you notice n is 2 when trying to find the derivative of x^2